### v1

@@ 3 A Method for Disentangling Generalization Mechanisms of GFlowNets
We consider 3 main experimental settings: (1) distilling (regressing to) flow functions, (2) memorization gaps in GFlowNets, and (3) offline and off-policy training regimes. Each experimental setup is built on a series of simplifying assumptions such that we can control for different moving parts and complexities that are present when training GFlowNets. We now present this experimental protocol, and then report and discuss our findings in §4.
@@ 3.1 Distilling (Regressing to) Flow Functions
Just as we can compute the distribution $p(x;\theta)$ over ${\mathcal{X}}$ exactly in this environment, we also compute edge flows $F(s\to s^{\prime})$ and $P_{F}(s^{\prime}|s)$ exactly. We refer to these as the true flow and forward policies, in opposition to the approximated $F(s\to s^{\prime};\theta)$ and $P_{F}(s^{\prime}|s;\theta)$. Note that we parameterize $F(s\to s^{\prime};\theta)$ and $P_{F}(s^{\prime}|s;\theta)$ as mappings from ${\mathcal{S}}\to{\mathbb{R}}^{|n(s)|}$ where $n(s)$ is the number of children of $s$.
In this set of experiments, we train DNNs by regressing to the true flow [^3] In this sense, we are “distilling” the true flows into flow functions parameterized by a DNN. $F(s\to s^{\prime})$ and forward policy $P_{F}(s^{\prime}|s)$. This removes the use of the GFlowNet training objective and the use of trajectories sampled from $P_{F}(s^{\prime}|s;\theta)$ as the training data distribution, drastically simplifying the training procedure. As such, we basically control for any non-ideal factors introduced by the GFlowNet training objectives, for example, possible shortcomings due to temporal credit assignment (Malkin et al., 2022).
We can now train models to regress to $\log F(s\to s^{\prime})$ or $\log P_{F}(s^{\prime}|s)$. These are both vectors, so we regress to each value independently and minimize the mean squared error. Note that the true $P_{F}(s^{\prime}|s)=\mathrm{softmax}(\log F(s\to s^{\prime}))$. Regressing to $\log F(s\to s^{\prime})$ is the same general problem as regressing to $\log P_{F}(s^{\prime}|s)$, but with one caveat. When learning $F(s\to s^{\prime};\theta)$, the model has to not only get the relative magnitudes of each logit right, but also the absolute magnitudes right. In contrast, when learning $P_{F}(s^{\prime}|s;\theta)$, because of the normalization within the softmax, the model only has to get the relative magnitude of the logits right. Because of this, we expect that learning $P_{F}(s^{\prime}|s)$ is easier, although learning flow magnitudes could help with generalization. For these experiments we set $P_{B}$ to be the uniform policy to get unique $F$ and $P_{F}$.
Using this setup we are able to asses how fundamentally difficult these flow functions are to learn. Furthermore, we can use our test set of unseen states. This allows us to asses generalization performance for learning $p(x)$ (which we are able to tractably compute) by directly computing distributional errors. Because of the use of a test set, this is in some sense more challenging than when training GFlowNets directly, i.e. online and on-policy, since standard training of GFlowNets allows the model to potentially explore the entire space and “see” the entire dataset [^4] Readers may notice that in this task, because we’ve computed the regression targets $F$ and $P_{F}$ exactly using the entire state space, some information of the test set is leaking into the targets. Since the purpose of this experiment is to asses how hard these functions fundamentally are to learn, we knowingly allow the model to cheat..
@@ 3.2 Memorization Gaps in GFlowNets
Problem Data Coupling
@@
Regress $R$ $s\!\perp\!\!\!\!\perp\tilde{R}(s)$
@@
Regress $P_{F}$ $(s,s^{\prime})\!\perp\!\!\!\!\perp P_{F}^{\mathrm{random}}$
Table 1: High level summary for the memorization gap experiments. Each row lists the setup of an individual experiment showing the corresponding data pair coupling and structure of learning problem.
@@
We would like to probe our hypothesis on learning true flows and their contribution to generalization in GFlowNets. Here, we take on the perspective of Zhang et al. (2017; 2021) and consider generalization as the act of not memorizing. In other words, we can assess whether a model is generalizing or not by measuring the gap (in training performance) between training it on structured data and training on random unstructured data. We use this notion of memorization gap to try and reason about learning structured flows and their implications on generalization.
We devise an experiment inspired from Zhang et al. (2017; 2021), where we train supervised models by, similarly to the previous section, regressing to $R(s)$ and $P_{F}(s^{\prime}|s)$, but with various degrees of “de-”structuring.
@@
Consider simply learning to predict $R(s)$ with $R(s;\theta)$. If instead of regressing from $s$ to $R(s)$ we shuffle the reward labels of each state $s$, we end up with new $(s,\tilde{R}(s))$ pairings. This induces independence between the data pairings, i.e. $s\!\perp\!\!\!\!\perp\tilde{R}(s)$. Intuitively, training a DNN with $s\!\perp\!\!\!\!\perp\tilde{R}(s)$ will force the DNN to memorize since we have removed any structured information that may exist in the problem. We compare the training curve gap between regressing to $R$ and $\tilde{R}$.
@@
Now consider regressing to $P_{F}(s^{\prime}|s)$; there are two ways to destructure this function. Recall that $P_{F}(s^{\prime}|s)\propto F(s^{\prime})$, and that $F(s)$ is a function of the reward of all its descendants as well as of the transition structure of the state space (the ways to get to those descendants). Let ${\mathcal{D}}(s)$ be the set of descendants of $s$. By shuffling $R$ into $\tilde{R}$, we thus remove the dependence between $s$ and $\tilde{R}(s_{d})\forall s_{d}\in{\mathcal{D}}(s)$, but keep the dependence between $s$ and how to get to ${\mathcal{D}}(s)$. The second way to de-structure $P_{F}$ is to simply regress to random logits, which we will denote $P_{F}^{\mathrm{random}}$.
To recap, we train models on both the coupled and permuted or random data. We consider 5 overall experiments (see Table 1), regressing to: $R(s)$ using paired data, $R(s)$ using shuffled data, $P_{F}(s^{\prime}|s)$ using paired data, $P_{F}(s^{\prime}|s)$ using shuffled data, and $P_{F}(s^{\prime}|s)$ using random logits $P_{F}^{\mathrm{random}}$ sampled (once before training) uniformly at random from a magnitude-preserving range. Table 1 summarizes the high level setup of each experiment and the corresponding data/flow structure of the problem.
@@
Through this setup, we can assess the memorization gap that occurs when training models on the respective settings and tasks. In particular, removing the structure of the data but maintaining flow structure in the learning problem allows us to assess the contribution that learning true flows has on generalization (i.e. not memorization) when training GFlowNets. Following from Hypothesis 2, if state flows have interesting structure and learning structured flows reduces the degree of memorization, perhaps this structure inherently acts as a mechanism for generalization in GFlowNets.
@@ 3.3 Offline and Off-policy Training Regimes
Lastly, we would like to investigate the effects of deviating from the self-induced training distributions of GFlowNets on generalization. To do this, we consider the setting of training GFlowNets offline and off-policy given a known dataset of final states ${\mathcal{X}}$. We sample $x\sim{\mathbb{P}}_{\mathcal{X}}$, where ${\mathbb{P}}_{\mathcal{X}}$ is some training distribution over ${\mathcal{X}}$. We consider different distributions for ${\mathbb{P}}_{\mathcal{X}}$ that resemble different practical approaches and techniques that are used for training GFlowNets (see Table 2). This setting controls for the effect of sampling from $P_{F}(s^{\prime}|s;\theta)$ in training GFlowNets.
@@
Given a terminal $x$, we sample a trajectory $\tau=(s_{0},\dots,s_{f})$ backwards using a uniform $P_{B}$ [^5] Investigating the use of different variants/distributions for $P_{B}$ is another interesting direction that may give insight into the effects of the backward policy on the training dynamics of GFlowNets (see Shen et al., 2023). We leave this for future work.. Hence, we term this as offline training, since we are avoiding sampling $x$ using $P_{F}(s^{\prime}|s;\theta)$. In this way, we are still able to use a standard GFlowNet objective, e.g. SubTB(1), to train $P_{F}(s^{\prime}|s;\theta)$ using the backward sampled trajectories. Note that this setup can remove the data non-stationarity that exists during normal training when using on-policy samples from $P(s^{\prime}|s;\theta)$.
@@
We also wish to assess to what degree deviating from $P_{F}(s^{\prime}|s;\theta)$ during training will affect learning $p(x)$. Understanding this may be useful to develop novel GFlowNet algorithms, considering that in practice it seems hard to train models that deviate from on-policy training (Rector-Brooks et al., 2023). To achieve this, we consider a policy interpolation experiment where we sample trajectories from $P_{\alpha}(s^{\prime}|s)=(1-\alpha)P_{F}(s^{\prime}|s;\theta)+\alpha P_{{\mathcal{U}}}(s^{\prime}|s)$, where $0\leq\alpha\leq 1$ and $P_{{\mathcal{U}}}(s^{\prime}|s)$ denotes a $P_{F}$ that goes to every terminal state with equal probability, i.e. such that $p(x)\propto 1$. We approximate $P_{{\mathcal{U}}}(s^{\prime}|s)$ using a policy pre-trained with a constant reward (thus approximating a uniform $p(x)$).
We then train $P_{F}(s^{\prime}|s;\theta)$ online instead of offline. Note that this is a slightly different setup from what we discussed previously since we re-introduce some of the effects of self-induced training distributions of $P_{F}(s^{\prime}|s;\theta)$. We do this because, as we will see, any degree of deviation from $P_{F}(s^{\prime}|s;\theta)$ during offline training (i.e. sampling $x\sim{\mathbb{P}}_{\mathcal{X}}$ and using $P_{B}$ to construct trajectories) appears to have some negative effect. We observe this even for values of $\alpha$ close to 0. We discuss this in further detail in section 4.4.
@@
This experimental setup helps us investigate the isolated effect of different data distributions ${\mathbb{P}}_{\mathcal{X}}$ on training dynamics and generalization by changing the data non-stationarity typically induced by $P(s^{\prime}|s;\theta)$. To observe the effects on training dynamics, we consider the setting of no test set (the model sees all of ${\mathcal{X}}$). In contrast, to observe the effects on generalization, we consider performance on unseen test set examples. This allows us to assess Hypothesis 1 by observing the effects of various deviations from the normal self-induced training distribution coming from $P_{F}(s^{\prime}|s;\theta)$.
@@ 4 Experiments
In this section, we report and discuss our observations and findings of the experimental methodology defined in §3. We conduct all experiments reported in this section over 3 random seeds. For training online and offline GFlowNets, we use SubTB(1) and a uniform $P_{B}$.
Figure 3: Training a model to distill edge flows and policies on 4 different tasks. It generally seems that (a) doing so recovers the intended distribution fairly well, and (b) modeling $P_{F}$ appears easier than modeling $F$, insofar as it recovers $p(x)$ better.
Figure 4: JS-divergence and MAE gaps between SubTB(1) trained GFlowNet model and $P_{F}$ distilled model. Distilling to $P_{F}$ generally appears to yield lower distributional error, with the exception of observing JS-divergence on the cliques task.
Figure 5: Training models distilled to $P_{F}$ and GFlowNet models trained online/on-policy for a range of monotonic skew values $\gamma$. Transforming the distribution of the reward to contain a larger proportion of high reward values generally does not yield significant influence on performance for approximating $p(x)$.
Figure 6: Memorization gap training curves for counting, neighbors, and cliques tasks. Maintaining flow structure in the learning problem (learning $P_{F}$ under shuffled $\tilde{R}$) induces a smaller memorization gap relative to the fully de-structured setting.
@@ 4.1 Distilling Flow Functions
We run the setup explained in §3.1. To evaluate the models, we look at the distributional errors on $p(x;\theta)$. We use a 90%-10% train-test split. We report results in Figure 3 and Figure 4.
First, we can see that learning $P_{F}$ and learning $F$ yield similar difficulty in the sense that the distributional errors are systematically close across tasks. Second, we see that the model is generalizing, getting sometimes even better scores than a model trained online on the entire states space (i.e. with no test set) as seen in Figure 4. The exception appears to be when evaluating the performance of the cliques task using the JS divergence metric, where the online GFlowNet yields marginally better performance than the distilled models. While expected, this second result is relevant, because it suggests that the difficulty of training a GFlowNet comes from both learning to model the flow functions themselves and performing temporal credit assignment.
@@
We confirm these observation in sequence and grid environments (see §E.1). We include further results in §D.1 where we assess how well GFlowNets assign probability mass to unvisited states in the graph generation tasks.
@@ 4.2 Reward complexity and transformations
Consider a monotonic transform ${\mathcal{H}}_{\gamma}:\mathbb{R}\rightarrow\mathbb{R}$ that maps $\log R^{\prime}(s)={\mathcal{H}}_{\gamma}(\log R(s)),\forall s\in{\mathcal{S}}$ such that $\log R^{\prime}(s)$ is skewed towards higher reward states as a function of some parameter $\gamma$ (see §C.3 for further details). We use the setup in §3.1 and also train online GFlowNets in the setting of skewed reward distributions transformed by ${\mathcal{H}}_{\gamma}$. Results are shown in Figure 5. We observe that as we increase $\gamma$ (increasing the skewness of the distribution of $\log R$ towards higher reward values), both distilled and online trained models are not significantly affected by the augmented reward distribution.
@@
Monotonic transformations of $R(x)$ do not have detrimental impact on generalization when learning $p(x)$ with GFlowNets.

### v2

@@ 3 A Method for Disentangling Generalization Mechanisms of GFlowNets
We consider 3 main experimental settings: (1) distilling (regressing to) flow functions, (2) memorization gaps in GFlowNets, and (3) offline and off-policy training regimes. Each experimental setup is built on a series of simplifying assumptions, allowing us control for different moving parts and complexities that are present when training GFlowNets. We now present this experimental protocol, and then report and discuss our findings in §4.
@@ 3.1 Distilling (Regressing to) Flow Functions
Just as we can compute the distribution $p(x;\theta)$ over ${\mathcal{X}}$ exactly in this environment, we also compute edge flows $F(s\to s^{\prime})$ and $P_{F}(s^{\prime}|s)$ exactly. We refer to these as the true flow and forward policies, in opposition to the approximated $F(s\to s^{\prime};\theta)$ and $P_{F}(s^{\prime}|s;\theta)$. Note that we parameterize $F(s\to s^{\prime};\theta)$ and $P_{F}(s^{\prime}|s;\theta)$ as mappings from ${\mathcal{S}}\to{\mathbb{R}}^{|n(s)|}$ with $n(s)$ the number of children of $s$.
In this set of experiments, we train DNNs by regressing to the true flow [^4] In this sense, we are “distilling” the true flows into flow functions parameterized by a DNN. $F(s\to s^{\prime})$ and forward policy $P_{F}(s^{\prime}|s)$. This removes the use of the GFlowNet training objective, and instead of the training distribution coming form trajectories sampled from $P_{F}(s^{\prime}|s;\theta)$ we sample $s$ from the training set, drastically simplifying the training procedure. As such, we basically control for any non-ideal factors within the GFlowNet training setup, such as shortcomings due to temporal credit assignment (Malkin et al., 2022).
To regress to $\log F(s\to s^{\prime})$ or $\log P_{F}(s^{\prime}|s)$, both vectors, we regress to each value independently and minimize the mean squared error. Note that the true $P_{F}(s^{\prime}|s)=\mathrm{softmax}(\log F(s\to s^{\prime}))$. Regressing to $\log F(s\to s^{\prime})$ is thus almost like regressing to $\log P_{F}(s^{\prime}|s)$, but the model has to get the absolute magnitudes of each logit right, not just the relative ones (the softmax in $P_{F}$ normalizes). Because of this, we expect that learning $P_{F}(s^{\prime}|s)$ is easier, although learning flow magnitudes could help with generalization. We set $P_{B}$ to be the uniform policy to get unique $F$ and $P_{F}$.
Using this setup we are able to assess how fundamentally difficult these flow functions are to learn. Furthermore, we can use our test set of unseen states. This allows us to assess generalization performance for learning $p(x)$ (which we are able to tractably compute) by directly computing distributional errors. Because of the use of a test set, this is in some sense more challenging than when training GFlowNets directly, i.e. online and on-policy, since standard training of GFlowNets allows the model to potentially explore the entire space and “see” the entire dataset [^5] Readers may notice that because we’ve computed the regression targets $F$ and $P_{F}$ exactly using the entire state space, some information of the test set is leaking into the targets. Since the purpose of this experiment is to assess how hard these functions fundamentally are to learn, we knowingly allow the model to cheat..
@@ 3.2 Memorization Gaps in GFlowNets
Table 3: High level summary for the memorization gap experiments. Each row lists an individual experiment showing the corresponding data pair coupling and structure of learning problem. These experiments consider models trained with distilled/regressed learning (not via online training using SubTB$(1)$). $\tilde{P}_{F}(s^{\prime}|s)$ denotes the policy logits derived from the shuffled rewards.
Learning
(Regressing to)
Data Coupling
@@
$R$ $s\!\perp\!\!\!\!\perp\tilde{R}(s)$
@@
$P_{F}$ $(s,s^{\prime})\!\perp\!\!\!\!\perp P_{F}^{\mathrm{random}}$
@@
We would like to probe our hypothesis on the contribution to generalization of learning flows in GFlowNets. We take the perspective of Zhang et al. (2017; 2021) and consider generalization as the act of not memorizing. In other words, we can assess whether a model is generalizing or not by measuring the gap (in training performance) between training it on structured data and training on random unstructured data. We use this notion of memorization gap to examine how learning flows with or without structure impacts generalization.
@@
We devise an experiment inspired from Zhang et al. (2017; 2021), where we train supervised models by regressing to $R(s)$ and $P_{F}(s^{\prime}|s)$ (using the framework described in the previous section), but with various degrees of “de-”structuring. Consider simply learning to predict $R(s)$ with $R(s;\theta)$. If instead of regressing from $s$ to $R(s)$ we shuffle the reward labels of each state $s$, we end up with new $(s,\tilde{R}(s))$ pairings. This induces independence between the data pairings, i.e. $s\!\perp\!\!\!\!\perp\tilde{R}(s)$. Intuitively, training a DNN with $s\!\perp\!\!\!\!\perp\tilde{R}(s)$ will force it to memorize, having removed structure in the mapping from $s$ to $R$.
@@
Now consider regressing to $P_{F}(s^{\prime}|s)$; there are two ways to destructure this function. Recall that $F(s)$ (and so implicitly $P_{F}(s\textquoteright|s)$) is a function of the reward of all its descendants as well as of the transition structure of the state space (the ways to get to those descendants). Let ${\mathcal{D}}(s)$ be the set of descendants of $s$. By shuffling $R$ into $\tilde{R}$, we thus remove the dependence between $s$ and $\tilde{R}(s_{d})\forall s_{d}\in{\mathcal{D}}(s)$, but keep the dependence between $s$ and how to get to ${\mathcal{D}}(s)$. The second way to de-structure $P_{F}$ is to simply regress to random logits, which we will denote $P_{F}^{\mathrm{random}}$.
To recap, we regress to: $R(s)$ using paired data, $R(s)$ using shuffled data, $P_{F}(s^{\prime}|s)$ using paired data, $P_{F}(s^{\prime}|s)$ using shuffled data, and $P_{F}(s^{\prime}|s)$ using random logits $P_{F}^{\mathrm{random}}$ from a magnitude-preserving range. This is summarized in Table 3.
@@
Through this setup, we can assess the memorization gap that occurs when training models. Destructuring the reward but maintaining flow structure in the learning problem allows us to assess the contribution that learning true flows has on generalization (i.e. not memorization) when training GFlowNets. Following from Hypothesis 2, if learning structured flows reduces the degree of memorization, perhaps flow prediction inherently acts as a mechanism for generalization in GFlowNets.
@@ 3.3 Offline and Off-policy Training Regimes
Lastly, we investigate the effects of deviating from the self-induced training distributions of GFlowNets on generalization. To do this, we consider the setting of training GFlowNets offline and off-policy given a known dataset of final states ${\mathcal{X}}$. We sample $x\sim{\mathbb{P}}_{\mathcal{X}}$, where ${\mathbb{P}}_{\mathcal{X}}$ is some training distribution over ${\mathcal{X}}$. We consider different distributions for ${\mathbb{P}}_{\mathcal{X}}$ that resemble different practical approaches and techniques that are used for training GFlowNets (see Table 4). This setting controls for the effect of sampling from $P_{F}(s^{\prime}|s;\theta)$ in training GFlowNets.
@@
Given a terminal $x$, we sample a trajectory $\tau=(s_{0},\dots,s_{f})$ backwards using a uniform $P_{B}$ [^6] Investigating the use of different $P_{B}$s is another interesting direction that may give insight into the effects of the backward policy on the training dynamics of GFlowNets (see Shen et al., 2023). We leave this for future work.. Hence, we term this as offline training, since we are avoiding sampling $x$ using $P_{F}(s^{\prime}|s;\theta)$.We use a standard GFlowNet objective, SubTB(1), to train $P_{F}(s^{\prime}|s;\theta)$ using those $\tau$. Note that this setup can remove the data non-stationarity that normally exists when using on-policy $P(s^{\prime}|s;\theta)$ samples.
@@
We also consider the online off-policy setting, to assess to what degree deviating from $P_{F}(s^{\prime}|s;\theta)$ during training will affect learning $p(x)$. Understanding this may be useful to develop novel GFlowNet algorithms, considering that in practice only few settings that deviate from on-policy training work (Rector-Brooks et al., 2023). To achieve this, we consider a policy interpolation experiment where we sample trajectories from $P_{\alpha}(s^{\prime}|s)=(1-\alpha)P_{F}(s^{\prime}|s;\theta)+\alpha P_{{\mathcal{U}}}(s^{\prime}|s)$, where $0\leq\alpha\leq 1$ and $P_{{\mathcal{U}}}(s^{\prime}|s)$ denotes a $P_{F}$ that goes to every terminal state with equal probability, i.e. such that $p(x)\propto 1$. We do this because, as we will see, any degree of deviation from $P_{F}(s^{\prime}|s;\theta)$ during offline training appears to have some negative effect. We observe this even for values of $\alpha$ close to 0. We discuss this in further detail in section 4.4.
@@
This experimental setup helps us investigate the isolated effect of different data distributions ${\mathbb{P}}_{\mathcal{X}}$ on training dynamics and generalization by changing the data distribution (and non-stationarity) typically induced by $P(s^{\prime}|s;\theta)$. We consider both settings with and without a test set. This allows us to assess Hypothesis 1 by observing the effects of various deviations from the normal self-induced training distribution coming from $P_{F}(s^{\prime}|s;\theta)$.
@@ 4 Experiments
In this section, we report and discuss our observations and findings of the experimental methodology defined in §3. We conduct all experiments reported in this section over 3 random seeds. For training online and offline GFlowNets, we use SubTB(1) (see §D.2) and a uniform $P_{B}$.
@@ 4.1 Distilling Flow Functions
Figure 2: (a) Training a model to distill edge flows and policies. (i) doing so recovers the intended distribution, (ii) modeling $P_{F}$ appears easier than modeling $F$. (b) JS-divergence and MAE gaps between SubTB(1) trained model and $P_{F}$ distilled model. Distilling to $P_{F}$ appears to yield lower distributional error, except for JS-divergence on the cliques task.
We run the setup explained in §3.1. To evaluate the models, we look at the distributional errors on $p(x;\theta)$. We use a 90%-10% train-test split. We report results in Figure 2(a) and Figure 2(b).
First, we can see that learning $P_{F}$ and learning $F$ yield similar difficulty in the sense that the distributional errors are systematically close. Second, we see that the model is generalizing, getting sometimes even better scores than a model trained online on the entire states space (i.e. with no test set) as seen in Figure 2(b). While this result should be obvious, it confirms that the difficulty of training a GFlowNet comes from both learning to model the flow functions themselves and performing temporal credit assignment.
@@
We confirm these observation in sequence and grid environments (see §E.1). Further results in §D.3 assess how well GFlowNets (and TB Malkin et al. (2022) and FM (Bengio et al., 2021)) generalize in rank in the graph tasks.
@@ 4.2 Reward complexity and transformations
Figure 3: (a) Training models distilled to $P_{F}$ and models trained online/on-policy for a range of $\gamma$; (b) similarly for a range of $\beta$. Transforming the distribution of the reward does not significantly affect the generalization difficulty in approximating $p(x)$.
Consider a monotonic transform ${\mathcal{H}}_{\gamma}:\mathbb{R}\rightarrow\mathbb{R}$ that maps $\log R^{\prime}(s)={\mathcal{H}}_{\gamma}(\log R(s)),\forall s\in{\mathcal{S}}$ such that $\log R^{\prime}(s)$ is skewed as a function of $\gamma$ (see §C.3 for further details). We use the setup in §3.1 and also train online GFlowNets in the setting of skewed reward distributions transformed by ${\mathcal{H}}_{\gamma}$ as well as tempered distributions via a $\beta$ parameter, $\log R^{\prime}(s)=\beta\log R(s)$. Results are in Figure 3(a); as we change $\gamma$ and $\beta$, both distilled and online models are not significantly affected.
@@
Monotonic transformations, such as skew and reward tempering, do not have detrimental impact on generalization when learning $p(x)$ with GFlowNets, for the range of hyperparameters considered.